Solve the given initial value problem using the method of Laplace transforms. y + 3y +2y = tu(t-3); y(0) = 0, y(0) = 1 ClProperties of Laplace Transforms L{f+g} = £{f} + L{g} L{cf} = CL{f} for any constant £{e atf(t)} (s) = L{f}(s-a) L{f}(s) = sTable of Laplace Transforms NWT šis>o e at ni t, n= 1,2,. sn + 1. S> 0 sin bt +62:> 0 cos bt 52+12,5> 0 n! (s-a)n +7 : Sa e
Answer #1

y + 3y+ 2y = tult-3) = (t-3+3) ult-3) = (t-3)ult-3)+ 36(t-3) Lly+ 3 Lly)+2167)=L(17-39Uct-3)) + 3 L(uct-3)) s?Lly) -8ycon-yPortal fractions froot (SHU)(8+2) = 6 + 2 ACS+2) +B(S+1)=1 S=+, A=1 and S=-2, -B=1 (5+1(8+2) = sti - str. SCF065+2) = 0 + 1 +